OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that has an x-intercept of -4 and a y-intercept of 3.

OpenStudy (anonymous):

Thank you!!

OpenStudy (midhun.madhu1987):

My Pleasure.....

OpenStudy (xapproachesinfinity):

your equation should look like this \(\Large \color{midnightblue}{y=mx+3}\) since -4 is x intercept you will have \(\Large \rm\color{midnightblue}{0=-4m+3}\) solving fo m \(\Large \rm\color{midnightblue}{m=3/4}\) so the equation becomes \(\Large \rm\color{midnightblue}{y=\frac{3}{4}x+3}\)